OpenStudy (anonymous):

The phone company Ringular has a monthly cellular plan where a customer pays a flat monthly fee and then a certain amount of money per minute used on the phone. If a customer uses 340 minutes, the monthly cost will be $132. If the customer uses 690 minutes, the monthly cost will be $237. A) Find an equation in the form y = m x + b, where x is the number of monthly minutes used and y is the total monthly of the Ringular plan. B)B) Use your equation to find the total monthly cost if 768 minutes are used. Me: So far I have the points (340,132) and (690,237)

OpenStudy (anonymous):

with the linear equation of y-132=3/7(x-340) i got 3/7 by using m=y2-y1/x2-x1 150/350= 3/7 but yet now im stuck